2. The domestic supply and demand curves for washing machines are as follows:

Supply: P= 2800+5Q Demand: P=4300-5Q

where P is the price in dollars and the Q is the quantity in millions. The U.S. is a small producer in the world washing machine market. Where the current price (which will not be affected by anything we do) is $ 3,000.

Congress is considering a tariff of $500.

A. Calculate and graph all points for the domestic market for the washing machine's price and quantity equilibrium.

B. Find the domestic quantity demanded and supplied of washing machines that will result if the price imposition of $3,000 is imposed.

Show on the graph.

Explain.

C. Find the domestic quantity demanded and supplied of washing machines that will result if the $500 tariff is imposed. Show on the graph.

Explain.

D. Compute government revenue from the tariff.
Answer #1

SS. .- P = 2800+50 i do = (P-2800/5 odd. P = 4300 - SQ Qd = (4300-P)/5 in the market when A) Equilibrium is achieved Qs = ddPrice (1) 4300 SS. aqui. old 8550 Pwt tariff 3500 3000 dd. 2800 0 40 Tho Tso 260 Quantity 160 is import at Pw import at Pot t